OpenStudy (anonymous):

y^2-12y+36=9 what are the exact solutions

Directrix (directrix):

y^2 -12y + 36 = 9 y^2 -12y + 36 is a perfect square trinomial. It factors as (y - 6) * (y - 6) or (y - 6)^2 so (y - 6)^2 = 9 (y - 6)^2 = 3^2 Taking square roots of both sides, y - 6 = 3 or y - 6 = -3 y = 3 +6 or y = -3 + 6 y = 9 or y = 3 ---> exact solutions We could have used the Quadratic Formula. Or, we could have factored after collecting all terms on one side. If you prefer one of those, I can help you get the same answers using one of those techniques. See what you think.

Directrix (directrix):

Question?

OpenStudy (anonymous):

Thanks so much but I am stuck in other areas right now